# Inductive Graph Representation Learning with Quantum Graph Neural Networks

### v2

@@ VI.2 Comparison Between GNNs and QGNNs
Each case is sampled randomly and independently, with 70% of the sample dataset used for training and 30% for testing. All experiments are conducted over 300 epochs and the results are presented in Figs. 6 and 7 for each case, showing the training and testing results of the GNN and QGNN frameworks. Table 2 reports the best test-set results, which we use to assess generalization performance. The corresponding training loss and score are also included.
@@
Despite the minimal design, these results highlight the potential of the proposed QGNN to process molecules of varying sizes and to offer advantages in generalization to unseen molecular structures. By benchmarking on a non-trivial dataset such as QM9, this study motivates further investigation of quantum convolutional graph architectures. Importantly, as is clearly shown in Figs. 6 and  7, the QGNN losses for both train and test set consistently plateaus during training across the two examined cases. Low model expressibility or the presence of suboptimal local minima in the loss landscape may be contributing factors that hinder the trainability of the quantum model [anschuetz_2021_critical, bittel_2021_training].
@@ VI.4 Scalability and absence of barren plateaus in QGNNs
Inspired by Ref. [pesah_physrevX_2021], we study the gradient scaling of our QGCN architecture under the same conditions. Specifically, as shown in the schematic representation of the circuit. 4, we consider two convolutional layer designs: an uncorrelated construction composed of independent local 2-designs $W$, and a correlated variant in which all $W$ operators within a layer share the same parameters. Finally, the loss function is defined by local measurements, ensuring its local nature due to the pooling structure of quantum convolutional models.

### v3

@@ VI.2 Comparison Between GNNs and QGNNs
Each case is sampled randomly and independently, with 70% of the sample dataset used for training and 30% for testing. All experiments are conducted over 300 epochs and the results are presented in Figs. 5 and 6 for each case, showing the training and testing results of the GNN and QGNN frameworks. Table 2 reports the best test-set results, which we use to assess generalization performance. The corresponding training loss and score are also included.
@@
Despite the minimal design, these results highlight the potential of the proposed QGNN to process molecules of varying sizes and to offer advantages in generalization to unseen molecular structures. By benchmarking on a non-trivial dataset such as QM9, this study motivates further investigation of quantum convolutional graph architectures. Importantly, as is clearly shown in Figs. 5 and  6, the QGNN losses for both train and test set consistently plateaus during training across the two examined cases. Low model expressibility or the presence of suboptimal local minima in the loss landscape may be contributing factors that hinder the trainability of the quantum model [anschuetz_2021_critical, bittel_2021_training].
@@ VI.4 Scalability and absence of barren plateaus in QGNNs
Inspired by Ref. [pesah_physrevX_2021], we study the gradient scaling of our QGCN architecture under the same conditions. Specifically, as shown in the schematic representation of the circuit. 3, we consider two convolutional layer designs: an uncorrelated construction composed of independent local 2-designs $W$, and a correlated variant in which all $W$ operators within a layer share the same parameters. Finally, the loss function is defined by local measurements, ensuring its local nature due to the pooling structure of quantum convolutional models.